OpenStudy (anonymous):

Solve for X: 3/x - 2/3 = 5/6

OpenStudy (anonymous):

3/x=5/6+2/3 3/x=5/6+4/6 3/x=9/6 now cross multiply 9x=18 x=2 :D

OpenStudy (anonymous):

Thank you(: